# Skull-to-Face: Anatomy-Guided 3D Facial Reconstruction and Editing

### v1

@@ Abstract
Deducing the 3D face from a skull is an essential but challenging task in forensic science and archaeology. Existing methods for automated facial reconstruction yield inaccurate results, suffering from the non-determinative nature of the problem that a skull with a sparse set of tissue depth cannot fully determine the skinned face. Additionally, their texture-less results require further post-processing stages to achieve a photo-realistic appearance. This paper proposes an end-to-end 3D face reconstruction and exploration tool, providing textured 3D faces for reference. With the help of state-of-the-art text-to-image diffusion models and image-based facial reconstruction techniques, we generate an initial reference 3D face, whose biological profile aligns with the given skull. We then adapt these initial faces to meet the statistical expectations of extruded anatomical landmarks on the skull through an optimization process. The joint statistical distribution of tissue depths is learned on a small set of anatomical landmarks on the skull. To support further adjustment, we propose an efficient face adaptation tool to assist users in tuning tissue depths, either globally or at local regions, while observing plausible visual feedback. Experiments conducted on a real skull-face dataset demonstrated the effectiveness of our proposed pipeline in terms of reconstruction accuracy, diversity, and stability. Our project page is https://xmlyqing00.github.io/skull-to-face-page/.
Keywords:
Skull-to-Face Reconstruction Human Face Tissue Depth Modeling Human Face Adaptation
1 Introduction
@@
Our proposed work builds upon two main observations. (1) Firstly, the biological profile, encompassing ancestry, age, gender, and stature of the body, can be estimated from skulls [47]. Different population groups exhibit distinct facial priors, and we argue that incorporating these is crucial for accurate facial reconstruction. We generate an initial guess of the 3D face with these attributes, whereas existing methods often overlook this key information. (2) Secondly, the combinations of tissue depths in real human faces follow intrinsic distributions that have not been well-studied. We introduce an additional control based on the distribution of tissue depth over facial reconstruction, providing an intuitive and effective tool for 3D face manipulation and exploration.
@@
Our approach includes an offline tissue depth distribution training stage and two online inference stages: semantic face model initialization, and anatomy-guided face adaptation. We also develop landmark-based face editing as an optional adjustment stage. (1) During the preprocessing stage, tissue depth distribution can be analyzed using available datasets of scanned skull and face pairs, or obtained from statistical data in forensic anthropology literature [43, 44, 48, 49]. (2) When processing a query skull, the key biological profile of the skull is used to generate a 2D realistic facial image, which is then reconstructed into a textured 3D face. Then, facial landmarks are generated, guided by the pre-trained tissue depth distribution and the geometry of the query skull. Finally, the face geometry is adjusted to align with these facial landmarks, resulting in the final skull-conformed face. (3) As an optional step, we provide a tool that allows users to intuitively adjust the face geometry by tuning a one-dimensional feature in the distribution space of tissue depths.
@@
The main contributions of this work are as follows: (1) We present a first approach that incorporates the biological profile into skull-to-face reconstruction. (2) By analyzing the distribution of tissue depth, we devise an additional control mechanism over the automatic reconstruction workflow, offering an intuitive and effective way to tune the face geometry. (3) Compared to existing methods, our method achieves state-of-the-art geometric accuracy qualitatively and quantitatively. We will release the codes and pre-trained models, making it the first publicly accessible approach in this field.
2 Related Work
2.1 Facial Reconstruction from Skull
Data-driven facial reconstruction.
@@ 2.1 Facial Reconstruction from Skull
Statistics-based methods for facial reconstruction employ high-dimensional vectors, typically representing faces through their 3D coordinates or feature points. These methods, using techniques such as PCA, learn the mapping between corresponding skulls and faces from training datasets [10, 39, 8, 12, 57, 33, 16, 22]. A new face is then generated as a linear combination of principal components. Other approaches [33, 14, 6, 39, 50, 30, 40] involve regression algorithms to predict faces from skulls. Zhang et al. [56] project 3D skulls to 2D images and leverage image-to-image style transferring networks [20, 58] for facial reconstruction. However, these data-driven methods require a significant number of joint skull-face scans, which are costly and difficult to acquire. The limited available datasets often result in overfitted networks and dataset bias.
Deformation-based facial reconstruction.
@@
2.2 Digital Face Synthesis
Constrained face editing.
@@ 3 Method
3.1 Pipeline Overview
Our skull-to-face generation pipeline has three components, as is shown in Fig.2. In the Tissue Depth Distribution (TDD) modeling, we analyze the tissue depth distribution in the training set and use the distribution to suggest a valid combination of tissue depths for the skull (Sec. 3.2). Then, when reconstructing a face from a query skull, we have two stages. In the Initial Face Generation stage, we utilize key biological profiles about the human body, i.e., gender, age, and ancestry to generate a 2D candidate portrait using Stable Diffusion XL [4] and lift it to a 3D face using the FLAME model [15] (Sec. 3.3). Finally, with the facial landmarks computed from the tissue depth model as the geometry constraints, we propose an Anatomy-guided Face Adaptation method to modify the 3D face according to facial landmark constraints. The details of the reconstructed face are preserved from the 2D candidate portrait image (Sec. 3.4).
3.2 Distribution of Tissue Depth
@@
(a) Three representative values of the first principal component are visualized overlaid on the tissue depth distribution.
(b) Adjusting the values of the first principal component provides easy control of tissue depths, and the face shapes change from thin to fat.
@@
3.2.1 TDD-Global Model
We analyze the tissue depth distribution using the training set of the skull-face dataset, which comprises $m$ samples, each with $n=78$ landmarks. It is important to note that the number of training samples $m$ in the publicly accessible skull-face dataset is limited ($m\leq 100$ in our experiments). Consequently, employing a categorized or neural-based model to fit the landmark distribution may be inappropriate. We therefore designed a simple yet effective way to model and control the tissue depths.
We first used principle component analysis (PCA) to project the joint distribution of tissue depths from an $n-$dimension space to a low-dimensional space. We observed that the first principal component predominantly governs the tissue depth distribution, accounting for $51.9\%$ of the variance. This finding significantly reduces the degree of freedom for the combination of 78 tissue depths. We projected joint tissue depths to the first component axis, denoted as $C$. The distribution of $C$ is visualized in Fig. 3(a). Furthermore, the effect of adjusting the values of the first principal component is illustrated In Fig. 3(b), where the directions of the tissues are depicted as blue sticks and the tissue depths in the training set are green spheres. A more comprehensive discussion of PCA components is included in the Supplementary.
When we monotonously sample values from the distribution $C$, the tissue depths change from short to long, and the corresponding face shape changes from thinner to rounder. To better visualize the relationship between face shape and the distribution $C$, we select mean values from the top, middle, and bottom $33\%$ of $C$ as representative combinations of tissue depths, and visualize them in Fig. 3(b). Note that while some related morphable face models also provide correlations between PCA components and face fullness [7], our model is the first model explicitly guided by anatomy, connecting facial skin to the underlying skull.
The above model enables us to modify face shapes based on the joint distributions of all the tissue depths on a skull. And we denote it as the TDD-Global model. Specifically, (a) we can use $C=0$ to sample the average tissue depths when we would like to generate average faces with moderate facial thickness. (b) we also can vary values in the first principal component to sample multiple plausible tissue depth combinations, thereby exploring different face shapes. The effectiveness of TDD-Global is validated in face interpolation and face shape modification experiments (see Sec. 4.3.1).
3.2.2 TDD-Regional Model
We integrate the above TDD-Global in our pipeline for automatic facial reconstruction. In addition, experts and investigators often require a manual tool to support local fine-tuning. The TDD-Global model alone does not allow for precise modification of specific facial landmarks without impacting other regions. On the other hand, modifying single tissue depth alone may result in a non-realistic face shape, because the tissue depths in a local region also adhere to a joint distribution. This feature is essential for detailed face editing. By enabling localized modification, the editing capability can be significantly enhanced.
@@ 3.2.2 TDD-Regional Model
According to the facial anatomy study [34, 37], facial musculature and neurovascular structures are located throughout faces and can be divided into several regions, each responding to different movements and expressions. Our TDD-Regional model can work effectively because it supports fine-tuning local face regions by adjusting regional tissue depths in accordance with relevant anatomical structures. We verified our TDD-Regional model by extensive experiments, as shown in Sec. 4.3.2 (e.g., Fig. 7) and the Supplementary video/document.
3.3 Initial Face Generation
When reconstructing a 3D face from a skull, in addition to geometric constraints, the biological profile derived from the skull can also aid in the reconstruction process. The biological attributes often include age, gender, and ancestry, which determine the basic characteristics of the generated face. In this first stage, we generate initial faces following the characteristics of the population group specified by these attributes.
We adopt Stable Diffusion XL [4] to generate high-resolution 2D portrait images from text prompts. We design four main dimensions of facial attributes: gender, ancestry, age, and face shape to generate various faces covering different populations. We also add additional prompts to ensure the diffusion model generates appropriate front faces.
We lift the synthesized portrait images to 3D faces by using DECA [15]. It encodes the image to latent space, consisting of shape, expression, pose, and albedo codes, then decodes the latent code to a 3D face with the FLAME representation. We transfer the texture and geometry details from the image to the 3D mesh. The generated faces conform to the biological profile and will be adapted to the final face by the geometric constraints of facial landmarks.
3.4 Anatomy-guided Face Adaptation
@@
The deviation between the aligned landmark constraints $\mathcal{\tilde{P}}^{f}$ and the landmarks from the generated face should be minimized:
@@
Since the corresponding landmarks on face $\mathcal{Q}^{f}$ are manually labeled, they are inevitably mismatched to the target facial landmarks $\mathcal{\tilde{P}}^{f}$. In this sense, $\mathcal{L}_{lmk}$ can be regarded as a regularization term. To improve the registration accuracy from the face geometry to the target landmarks, we project the target facial landmark constraints $\mathcal{\tilde{P}}^{f}$ to the reconstructed face as $\mathcal{\tilde{P}}^{pf}$. Then we minimize the Euclidean distance between $\mathcal{\tilde{P}}^{f}$ and $\mathcal{\tilde{P}}^{pf}$ as a data term to constrain the face shape,
$\mathcal{L}_{proj}=\sum_{p\in\mathcal{\tilde{P}}^{pf}}\underset{v\in F}{\min}\|v-p\|_{2}^{2},$ (4)
@@
For any landmark on the left side $\mathbf{q}^{f}_{i}\in\mathcal{Q}^{f}_{L}$, we denote $\mathbf{q}^{pf}_{i}\in\mathcal{Q}^{f}_{R}$ as the corresponding landmark on the right side. We also denote the reflected point across the plane $\Pi$ as $\mathbf{q}^{rf}_{i}$. The symmetry loss function is defined as the Euclidean distance between the actual midpoint $\mathbf{q}^{M}$ and the expected midpoint $\mathbf{q}^{M^{\prime}}$,
$\mathbf{q}^{M}=\big{(}\mathbf{q}^{f}_{i}+$ $\mathbf{q}^{pf}_{i}\big{)}/2,\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \mathbf{q}^{M^{\prime}}=(\mathbf{q}^{f}_{i}+\mathbf{q}^{rf}_{i})/2,$ (5)
@@
4 Experiment
4.1 Dataset and Error Metric
We adopted a CT-scan dataset, named Skull100, comprising 100 pairs of 3D skulls and faces. These pairs were reconstructed (using [31]) from whole-head CT scans. Due to the privacy issue, this dataset doesn’t collect face images or appearance data. Details of dataset collection are explained in the supplementary. Skull100 is used as the ground truth to evaluate faces reconstructed from their corresponding skulls. We divide the dataset into two subsets: 50 pairs for training and 50 for testing. This large test set size is chosen to ensure the statistical significance of our evaluation results. We conducted the K-Fold cross-validation to evaluate the distribution bias (Sec. 4.4.2).

### v2

@@ Abstract
Deducing the 3D face from a skull is a challenging task in forensic science and archaeology. This paper proposes an end-to-end 3D face reconstruction pipeline and an exploration method that can conveniently create textured, realistic faces that match the given skull. To this end, we propose a tissue-guided face creation and adaptation scheme. With the help of the state-of-the-art text-to-image diffusion model and parametric face model, we first generate an initial reference 3D face, whose biological profile aligns with the given skull. Then, with the help of tissue thickness distribution, we modify these initial faces to match the skull through a latent optimization process. The joint distribution of tissue thickness is learned on a set of skull landmarks using a collection of scanned skull-face pairs. We also develop an efficient face adaptation tool to allow users to interactively adjust tissue thickness either globally or at local regions to explore different plausible faces. Experiments conducted on a real skull-face dataset demonstrated the effectiveness of our proposed pipeline in terms of reconstruction accuracy, diversity, and stability. Our project page is https://xmlyqing00.github.io/skull-to-face-page/.
Index Terms:
Skull-to-Face reconstruction, human face tissue depth modeling, human face adaptation.
I Introduction
@@
Our proposed work builds upon two main observations. (1) Firstly, the biological profile, encompassing ancestry, age, gender, and stature of the body, can be estimated from skulls [23]. Different population groups exhibit distinct facial priors, and we argue that incorporating these is crucial for accurate facial reconstruction. We generate an initial guess of the 3D face with these attributes, whereas existing methods often overlook this key information. (2) Secondly, the combinations of tissue depths in real human faces follow intrinsic distributions that have not been well-studied. We proposed a 3D face representation based on the skull and the tissue depth distribution models. This representation can precisely control the geometry shape on the 3D face. We introduce an additional control based on the distribution of tissue depth over facial reconstruction, providing an intuitive and effective tool for 3D face manipulation and exploration.
@@
Our approach includes an offline training stage for tissue depth distribution and two-stage online inference for face reconstruction: semantic face model initialization, and anatomy-guided face adaptation. We also develop landmark-based face editing as an optional adjustment stage. (1) During the preprocessing stage, tissue depth distribution can be analyzed using available datasets of scanned skull and face pairs, or obtained from statistical data in forensic anthropology literature [19, 20, 21, 22, 24, 25]. (2) When processing a query skull, the key biological profile of the skull is used to generate a 2D realistic facial image, which is then reconstructed into a textured 3D face. Then, facial landmarks are generated, guided by the pre-trained tissue depth distribution and the geometry of the query skull. Finally, the face geometry is adjusted to align with these facial landmarks, resulting in the final skull-conformed face. (3) As an optional step, we provide a tool that allows users to intuitively adjust the face geometry by tuning a one-dimensional feature in the distribution space of tissue depths.
The main contributions of this work are as follows:
  • (1)
    We present a novel face generation and adaptation pipeline that can effectively reconstruct a face from a given skull. This pipeline incorporates a new 3D face representation where faces are defined by learned tissue depth (thickness) distributions. This pipeline supports (a) generation of faces that accurately follow geometry constraint from the skull, and (b) intuitive exploration of plausible faces that match this skull.
  • (2)
    We leverage foundation models to create a bio-profile-aligned initial face. By learning tissue depth distributions, we design an effective face adaption module to modify the initial face to match the query skull.
  • (3)
    The learned tissue depth distribution also allows us to build a control mechanism that offers an intuitive and effective way to explore different plausible faces.
  • (4)
@@
    Compared to existing methods, our method achieves state-of-the-art geometric accuracy qualitatively and quantitatively. We will release the codes and pre-trained models, making it the first publicly accessible approach in this field.
II Related Work
II-A Facial Reconstruction from Skull
@@
Statistics-based methods for facial reconstruction employ high-dimensional vectors, typically representing faces through their 3D coordinates or feature points. These methods, using techniques such as PCA, learn the mapping between corresponding skulls and faces from training datasets [11, 12, 13, 14, 26, 27, 15, 28]. A new face is then generated as a linear combination of principal components. Other approaches [27, 29, 30, 12, 31, 32, 33] involve regression algorithms to predict faces from skulls. Zhang et al. [10] project 3D skulls to 2D images and leverage image-to-image style transferring networks [34, 35] for facial reconstruction. SCULPTOR [33] learns a joint distribution of skull and face from the annotated dataset, but it’s hard to integrate it into our system because only a few sample data and partial codes are publicly available. However, these data-driven methods require a significant number of joint skull-face scans, which are costly and difficult to acquire. The limited available datasets often result in overfitted networks and dataset bias.
@@
II-B Physics-based Face Modeling
Recent physics-based approaches for facial animation can be categorized into two types: jaw-driven and muscle-driven. Zoss et al. [37] presents a method to learn the non-linear mapping from the skin deformation to the underlying jaw motion on an annotated dataset. This mapping can be re-targeted to new subjects to guide the facial animation. Sifakis et al. citesifakis2005automatic built a model of facial musculature, passive tissue, and underlying skeletal structure using volumetric data acquired from a living male subject. They used the finite element model to drive the face animation by muscle activation. Ichim et al. [38] model the physical interaction of passive flesh, active muscles, and rigid bone structures by minimizing a set of non-linear potential energies. The learned muscle activation model leads to a robust reproduction of complex facial articulations. Kadleček and Kavan [39] built a volumetric model from magnetic resonance images of a neutral facial expression. They solved an inverse physics problem to learn the mechanical properties of the face from the training data. Our method provides another solution that uses tissue depths to control the shape of the human face.
II-C Digital Face Synthesis
@@ III Method
III-A Pipeline Overview
Our skull-to-face generation pipeline has three components, as is shown in Fig.2. In the Tissue Depth Distribution (TDD) modeling, we analyze the tissue depth distribution in the training set and use the distribution to suggest a valid combination of tissue depths for the skull (Sec. III-B). Then, when reconstructing a face from a query skull, we have two stages. In the Initial Face Generation stage, we utilize key biological profiles about the human body, i.e., gender, age, and ancestry to generate a 2D candidate portrait using Stable Diffusion XL [47] and lift it to a 3D face using the FLAME model [51] (Sec. III-C). Finally, with the facial landmarks computed from the tissue depth model as the geometry constraints, we propose an Anatomy-guided Face Adaptation method to modify the 3D face according to facial landmark constraints. The details of the reconstructed face are preserved from the 2D candidate portrait image (Sec. III-D).
III-B Distribution of Tissue Depth
@@
Medical and forensic studies have collected several soft tissue depth databases [24, 25] that serve as valuable references for tissue depth. These databases enable inference of facial tissue depth based on factors such as BMI, sex, age, and regional population. This option allows users to retrieve corresponding tissue depths from the database when available. However, since a comprehensive database that includes all population groups is unfeasible, we provide a generalized learning model to represent the typical distribution of human tissue depths.
III-B1 TDD-Global Model
We analyze the tissue depth distribution using the training set of the skull-face dataset, which comprises $m$ samples, each with $n=78$ landmarks. It is important to note that the number of training samples $m$ in the publicly accessible skull-face dataset is limited ($m\leq 100$ in our experiments). Consequently, employing a categorized or neural-based model to fit the landmark distribution may be inappropriate. We therefore designed a simpler model to encode the tissue depth distribution.
We first used principle component analysis (PCA) to project the joint distribution of tissue depths from an $n-$dimension space to a low-dimensional space. We observed that the first principal component predominantly governs the tissue depth distribution, accounting for $51.9\%$ of the variance. This finding significantly reduces the degree of freedom for the combination of 78 tissue depths. Consequently, we projected joint tissue depths to the first component axis, denoted as $C$. The distribution of $C$ is visualized in Fig. 3(a). Furthermore, the effect of adjusting the values of the first principal component is illustrated In Fig. 3(b), where the directions of the tissues are depicted as blue sticks and the tissue depths in the training set are green spheres. A more comprehensive discussion of PCA components is included in the Supplementary.
When we monotonously sample values from the distribution $C$, the tissue depths change from short to long, and the corresponding face shape changes from thinner to rounder. To better visualize the relationship between face shape and the distribution $C$, we select mean values from the top, middle, and bottom $33\%$ of $C$ as representative combinations of tissue depths, and visualize them in Fig. 3(b). Note that while some related morphable face models also provide correlations between PCA components and face fullness [55], our model is the first model explicitly guided by tissue depth distribution, directly connecting facial skin to the underlying skull.
The above model enables us to modify face shapes based on the joint distributions of all the tissue depths on a skull. And we denote it as the TDD-Global model. The TDD-Global model is essential because each value of $C$ represents a valid combination of tissue depths, allowing for easy adjustments to the overall tissue depths based on geometric constraints. Specifically, (a) we can use $C=0$ to sample the average tissue depths when we would like to generate average faces with moderate facial thickness. (b) We can generate multiple facial reconstructions with different values of $C$ to explore different face shapes and allow the user to select the best fit. (c) We can select $C$ based on available statistical tissue depth data, such as BMI or fatness level, to align more closely with the specific characteristics of the query skull. The effectiveness of TDD-Global is validated in face interpolation and face shape modification experiments (see Sec. IV-C1).
(a) Three representative values of the first principal component
are visualized overlaid on the tissue depth distribution.
(b) Adjusting the values of the first principal component
provides easy control of tissue depths, and the face shapes
change from thin to fat.
@@ III-B1 TDD-Global Model
III-B2 TDD-Regional Model
We integrate the above TDD-Global in our pipeline for automatic facial reconstruction. The correlation between regions is encoded in the global control model. In addition, experts and investigators often require a manual and interactive tool to support local fine-tuning. The TDD-Global model alone does not allow for precise modification of specific facial landmarks without impacting other regions. On the other hand, modifying single tissue depth alone may result in a non-realistic face shape, because the tissue depths in a local region also adhere to a joint distribution. This feature is essential for detailed face editing. Hence, we introduced the TDD-Regional Model that offers users additional control for fine-tuning details. By enabling localized modification, we intentionally remove interdependencies between regions, and the editing capability can be significantly enhanced.
@@ III-B2 TDD-Regional Model
According to the facial anatomy study [56, 57], facial musculature and neurovascular structures are located throughout faces and can be divided into several regions, each responding to different movements and expressions. Our TDD-Regional model can work effectively because it supports fine-tuning local face regions by adjusting regional tissue depths in accordance with relevant anatomical structures. We provide TDD-Regional as an additional control option for users, alongside our automatic TDD-Global pipeline. Our focus is on decomposing tissue depths for visualization purposes, rather than analysis. We verified our TDD-Regional model by extensive experiments, as shown in Sec. IV-C2 (e.g., Fig. 7) and the Supplementary video/document.
III-C Initial Face Generation
When reconstructing a 3D face from a skull, in addition to geometric constraints, the biological profile derived from the skull can also aid in the reconstruction process. The biological attributes often include age, gender, and ancestry, which determine the basic characteristics of the generated face. We provide users with two options to generate an initial face through text prompt depending on how much information of the query skull is available. By default, when the user doesn’t have much information of the skull, they can use the provided template prompt with attributes to synthesize the initial face. Each attribute has several potential choices for the user to select. In the other situation, when rich information of the skull is available, the user can add specific descriptions or disable some descriptions in the template to adjust the text prompt to adapt to the biological profile of the skull. This could generate a more detailed initial face as needed. In this first stage, we generate initial faces following the characteristics of the population group specified by these attributes.
We adopt Stable Diffusion XL [47] to generate high-resolution 2D portrait images from text prompts. We design four main dimensions of facial attributes: gender, ancestry, age, and face shape to generate various faces covering different populations. We also add additional prompts to ensure the diffusion model generates appropriate front faces. In addition to text prompts, users can also use other image generator such as ControlNet [58] to take edges, scribbles, and sketches as input to synthesize the initial face.
We lift the synthesized portrait images to 3D faces by using DECA [51]. It encodes the image to latent space, consisting of shape, expression, pose, and albedo codes, then decodes the latent code to a 3D face with the FLAME representation. We transfer the texture and geometry details from the image to the 3D mesh. The face initially generated by Stable Diffusion based on the biological profile serves as a preliminary estimate. It is then refined using an optimization-based method to align with the skull. The final 3D face shape is determined by the tissue depth measurements and skull structure, ensuring strict adherence to geometric constraints.
III-D Anatomy-guided Face Adaptation
@@
The deviation between the aligned landmark constraints $\mathcal{\tilde{P}}^{f}$ and the landmarks $\mathcal{Q}^{f}$ from the generated face should be minimized:
@@
Due to the limited resolution of the face mesh, the landmarks on the face $\mathcal{Q}^{f}$ may not perfectly align with the target facial landmarks $\mathcal{\tilde{P}}^{f}$. To address this, we add a constraint that brings the adapted face mesh as close as possible to the facial landmarks. We calculate the shortest distance from each target landmark to the adapted mesh and minimize this distance as a projection loss $\mathcal{L}_{proj}$.
$\mathcal{L}_{proj}=\sum_{p\in\mathcal{\tilde{P}}^{f}}\underset{v\in F}{\min}\|v-p\|_{2}^{2},$ (4)
@@
For any landmark on the left side $\mathbf{q}^{f}_{i}\in\mathcal{Q}^{f}_{L}$, we denote $\mathbf{q}^{pf}_{i}\in\mathcal{Q}^{f}_{R}$ as the corresponding landmark on the right side. We also denote the reflected point across the plane $\Pi$ as $\mathbf{q}^{rf}_{i}$. The symmetry loss function is defined as the Euclidean distance between the actual and expected midpoints $\mathbf{q}^{M}$, $\mathbf{q}^{M^{\prime}}$,
$\mathbf{q}^{M}=\big{(}\mathbf{q}^{f}_{i}+$ $\mathbf{q}^{pf}_{i}\big{)}/2,~{}~{}~{}\mathbf{q}^{M^{\prime}}=(\mathbf{q}^{f}_{i}+\mathbf{q}^{rf}_{i})/2,$ (5)
@@
IV Experiment
IV-A Dataset and Error Metric
We adopted a CT-scan dataset, named Skull100, comprising 100 pairs of 3D skulls and faces. These pairs were reconstructed (using [61]) from whole-head CT scans. Due to the privacy issue, this dataset doesn’t collect face images or appearance data. Details of dataset collection are explained in the supplementary. Skull100 is used as the ground truth to evaluate faces reconstructed from their corresponding skulls. We divide the dataset into two subsets: 50 pairs for training and 50 for testing. This large test set size is chosen to ensure the statistical significance of our evaluation results. We conducted the K-Fold cross-validation to evaluate the distribution bias (Sec. IV-D2).